# AMUN: Adversarial Machine UNlearning

### v2

@@ 1 Introduction
The goal of machine unlearning is to remove the influence of a subset of the training dataset for a model that has been trained on that dataset (Vatter et al., 2023). The necessity for these methods has been determined by privacy regulations such as the European Union’s General Data Protection Act and the California Consumer Privacy Act. Despite early efforts on proposing “exact” solutions to this problem (Cao and Yang, 2015; Bourtoule et al., 2021), the community has favored “approximate” solutions due to their ability to preserve the original model’s accuracy while being more computationally efficient (Chen et al., 2023b; Liu et al., 2024; Fan et al., 2024).
@@
As we will show in § 6, AMUN outperforms prior state-of-the-art (SOTA) unlearning methods (Fan et al., 2024) in unlearning random subsets of the training data from a trained classification model and closes the gap with the retrained models, even when there is no access to the samples in $\mathcal{D}-\mathcal{D_{\text{F}}}$ during the unlearning procedure.
@@ 1.1 Related Work
Early works in machine unlearning focused on exact solutions Cao and Yang (2015); Bourtoule et al. (2021); those ideas were adapted to unlearning in other domains such as graph neural networks Chen et al. (2022b) and recommendation systems Chen et al. (2022a). The extensive computational cost and utility loss resulted in the design of approximate methods. An example is the work of Ginart et al. (2019), who provide a definition of unlearning based on differential privacy. Works that followed sought solutions to satisfy those probabilistic guarantees (Ginart et al., 2019; Gupta et al., 2021; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021). However, the methods that satisfy these guarantees were only applied to simple models, such as $k$-means Ginart et al. (2019) , linear and logistic regression Guo et al. (2019); Izzo et al. (2021), convex-optimization problems Neel et al. (2021), or graph neural networks with no non-linearities Chien et al. (2022). Additional research was carried out to design more scalable approximate methods, those that can be applied to the models that are used in practice, including large neural networks Golatkar et al. (2020); Warnecke et al. (2021); Izzo et al. (2021); Thudi et al. (2022); Chen et al. (2023b); Liu et al. (2024); Fan et al. (2024). However, these approximate methods do not come with theoretical guarantees; their effectiveness are evaluated using membership inference attacks (MIAs). MIAs aim to determine whether a specific data sample was used in the training set of a trained model (Shokri et al., 2017; Yeom et al., 2018; Song et al., 2019; Hu et al., 2022; Carlini et al., 2022; Zarifzadeh et al., 2024), and is a common evaluation metric (Liu et al., 2024; Fan et al., 2024). s For further discussion on related works see Appendix B.
@@ 3.1 A Guiding Observation
Results: Figure 3 in Appendix E.1, shows the confidence scores (see § E.1 for details) for a ResNet-18 He et al. (2016) model that has been retrained on $\mathcal{D}-\mathcal{D_{\text{F}}}$, where $\mathcal{D}$ is the training set of CIFAR-10 Alex (2009) and the size of $\mathcal{D_{\text{F}}}$ (randomly chosen from $\mathcal{D}$) is $10\%$ and $50\%$ of the size of $\mathcal{D}$ (the first and second sub-figures, respectively).
@@
This basic observation has either been overlooked by the prior research on approximate machine unlearning or has been treated incorrectly. To make the unlearned models more similar to $\mathcal{F_{\text{R}}}$, prior methods have focused on degrading the model’s performance on $\mathcal{D_{\text{F}}}$ directly by either (a) some variation of fine-tuning on $\mathcal{D_{\text{R}}}$ (Warnecke et al., 2021; Liu et al., 2024), (b) choosing wrong labels for samples in $\mathcal{D_{\text{F}}}$ and fine-tuning the model (Golatkar et al., 2020; Chen et al., 2023b; Fan et al., 2024), or (c) directly maximizing the loss with respect to the samples in $\mathcal{D_{\text{F}}}$ (Thudi et al., 2022). Using the wrong labels for the samples of $\mathcal{D_{\text{F}}}$ or maximizing the loss on them make these methods very unstable and prone to catastrophic forgetting (Zhang et al., 2024a) because these samples belong to the correct distribution of the data and we cannot force a model to perform wrongly on a portion of the dataset while preserving it’s test accuracy. For these methods, it is important to use a small enough learning rate along with early stopping to prevent compromising the model’s performance while seeking worse prediction confidence values on the samples in $\mathcal{D_{\text{F}}}$. Also, most of these methods require access to the set of remaining samples to use it for preventing a total loss of the model’s performance (e.g., by continuing to optimize the model on $\mathcal{D_{\text{R}}}$(Golatkar et al., 2020; Liu et al., 2024).
@@ 3.2 Fine-tuning on Adversarial Examples
Now, what happens if we insert one adversarial example $(x_{adv},y_{adv})$ that corresponds to the sample $(x,y)$ into $\mathcal{D}$ and make an augmented dataset $\mathcal{D}^{\prime}$ for fine-tuning? Even before fine-tuning starts, the model makes the correct prediction on that (adversarial) example (by predicting the wrong label $y_{adv}$!), but its confidence might not be as high as the samples in $\mathcal{D}$, on which the model has been trained on. Proceeding with fine-tuning of the model on the augmented dataset increases its confidence on $x_{adv}$ while making the same wrong prediction $y_{adv}$. However, this fine-tuning does not change the model’s performance because the newly added sample $(x_{adv},y_{adv})$ does not contradict the distribution learned by the model. Since $(x,y)\in\mathcal{D}^{\prime}$, and $x$ and $x_{adv}$ are very close to one another (e.g., very similar images) while having different labels, optimizing the model has to change its decision boundary in that region of the input space to reach small loss for both of these samples. As a result of this balance, the model tends to decrease its confidence on the original sample compared to the model that was solely trained on $\mathcal{D}$ because there was no opposing components for its optimization on $\mathcal{D}$. Note that $\|x-x_{adv}\|\leq\epsilon$, where $\epsilon$ is often much smaller than the distance of any pairs of samples in $\mathcal{D}$. This helps to localize this change in the decision boundary during fine-tuning, and prevent changes to models’ behavior in other regions of the input space (Liang et al., 2023). In the following we elaborate on our empirical observations that verify these changes.
@@ 4 Adversarial Machine UNlearning (AMUN)
Formally, AMUN uses Algorithm 1 to find an adversarial example for any sample in $(x,y)\in\mathcal{D_{\text{F}}}$. This algorithm uses a given untargeted adversarial algorithm $\mathcal{A}_{\mathcal{F}}$, that finds the solution to Definition 2.1, for finding an adversarial example $x_{adv}$. To make sure $\epsilon$ is as small as possible, Algorithm 1 starts with a small $\epsilon$ and runs the attack $\mathcal{A}_{\mathcal{F}}$; if an adversarial algorithm is not found within that radius, it runs $\mathcal{A}_{\mathcal{F}}$ with a larger $\epsilon$. It continues to perform $\mathcal{A}_{\mathcal{F}}$ with incrementally increased $\epsilon$ values until it finds an adversarial example; it then adds it to $\mathcal{D_{\text{A}}}$. The algorithm stops once it finds adversarial examples for all the samples in $\mathcal{D_{\text{F}}}$.
The reason behind minimizing the distance of $\epsilon$ for each sample is to localize the changes to the decision boundary of the model as much as possible; this prevents changing the model’s behavior on other parts of the input space. For our experiments, we use PGD-50 (Madry, 2017) with $\ell_{2}$ norm bound as $\mathcal{A}_{\mathcal{F}}$. We set the step size of the gradient ascent in the attack to $0.1\times\epsilon$, which changes with the $\epsilon$ value. More details regarding the implementations of AMUN and prior unlearning methods and tuning their hyper-parameters can be found in Appendix C. Also, in Appendix E.5, we will show how using weaker attacks, such as Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014), might lead to lower performance of AMUN.
@@ 4.1 Influencing Factors
In this section, we derive an upper-bound on the $2$-norm of the difference of the parameters of the unlearned model and the retrained model (which are gold-standard for unlearning) that illuminated the influencing factors in the effectiveness of AMUN. To prove this theorem, we make assumptions that are common in the certified unlearning literature. The proof is given in Appendix A.
@@
Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=\{1,\dots,N\}}$ be a dataset of $N$ samples and without loss of generality let $(x_{n},y_{n})$ (henceforth represented as $(x,y)$ for brevity) be the sample that needs to be forgotten and $(x^{\prime},y^{\prime})$ be its corresponding adversarial example used by AMUN such that $\|x-x^{\prime}\|_{2}=\delta$. Let ^$\hat{\mathcal{R}}(w)$ represent the (unnormalized) empirical loss on $\mathcal{D}^{\prime}=\mathcal{D}\cup\{(x^{\prime},y^{\prime})\}$ for a model $f$ that is parameterized with $w$. We assume that $f$ is $L$-Lipschitz with respect to the inputs and $\hat{\mathcal{R}}$ is $\beta$-smooth and convex with respect to the parameters. Let $\theta_{o}$ represent the parameters corresponding to the model originally trained on $\mathcal{D}$ and $\theta_{u}$ be the parameters derived when the model is trained on $\mathcal{D}-\{(x,y)\}$. We also assume that both the original and retrained models achieve near-00 loss on their training sets. After AMUN performs fine-tuning on $\mathcal{D}^{\prime}$ using one step of gradient descent with a learning rate of $\frac{1}{\beta}$ to derive parameters $\theta^{\prime}$, we get the following upper-bound for the distance of the unlearned model and the model retrained on $\mathcal{D}-\{(x,y)\}$ (gold standard of unlearning):
@@
Note that the first two implications rely on the strength of the adversarial example in addition to being close to the original sample. The second bullet, which relies on the transferability of adversarial examples, has been shown to improve as the Lipschitz constant decreases (Ebrahimpour-Boroojeny et al., 2024). The third bullet point is a natural implication which also holds for other unlearning methods that rely on the fine-tuning of the model. The fourth bullet point is not relevant to the unlearning method and instead relies on the fact that the retrained model should have good generalizability to unseen samples; it implies that as the size of $\mathcal{D_{\text{F}}}$ increases (i.e., $|\mathcal{D_{\text{R}}}|$ decreases) and the performance of the retrained model decreases, the effectiveness of the unlearning model also decreases. This is also intuitively expected in the unlearning process. Hence, the proved theorem also justifies our earlier intuitions about the need for good quality adversarial examples that are as close as possible to the original samples (which is the goal of Algorithm 1).
@@ 5.1 Baseline Methods
We compare AMUN with FT (Warnecke et al., 2021), RL (Golatkar et al., 2020), GA (Thudi et al., 2022), BS (Chen et al., 2023b), $l_{1}$-Sparse (Liu et al., 2024), and SalUn (Fan et al., 2024). We also combine the weight saliency idea for masking the model parameters to limit the changes to the parameters during fine-tuning with AMUN and present its results as AMUN+SalUn (see Appendix D for more details). We use the same hyper-parameter tuning reported by prior works. For further details, see Appendix C.
@@ 5.2 Evaluation Metrics
The metic used by recent works in unlearning to evaluate the unlearning methods (Liu et al., 2024; Fan et al., 2024) considers the models retrained on $\mathcal{D_{\text{R}}}$ as the goal standard for comparison. They compute the following four values for both the retrained models and the models unlearned using approximate methods:
@@
The MIAs used in the recent unlearning methods by Liu et al. (2024); Fan et al. (2024) are based on the methods introduced by Yeom et al. (2018); Song et al. (2019). Although these MIAs have been useful for basic comparisons, recent SOTA MIAs significantly outperform their earlier counterparts, albeit with an increase in complexity and computation cost. To perform a comprehensive comparison of the effectiveness of the unlearning methods, we utilized a SOTA MIA called RMIA (Zarifzadeh et al., 2024), in addition to using the MIAs from prior works. In RMIA, the area under the ROC curve (AUC) of the MIA scores for predicting the training samples from the unseen samples is reported. Recall that in machine unlearning, the samples are split to three sets: $\mathcal{D_{\text{R}}}$, $\mathcal{D_{\text{F}}}$, and $\mathcal{D_{\text{T}}}$. For an unlearning method to be effective, as discussed in § 3.1, we expect the AUC of RMIA for distinguishing the samples in $\mathcal{D_{\text{F}}}$ from the ones in $\mathcal{D_{\text{T}}}$ to be the same as random guessing ($50\%$ assuming balanced data). As shown in Table 1, this expectation holds for the models retrained on $\mathcal{D_{\text{R}}}$.
@@ 5.3 Unlearning Settings
Another important factor missing in the comparisons of the unlearning methods in prior works is the possibility of access to $\mathcal{D_{\text{R}}}$. So, for our experiments we consider two settings, one with access to $\mathcal{D_{\text{R}}}$ and one with access to only $\mathcal{D_{\text{F}}}$. We adapt each of the unlearning methods to both of these settings, and perform the comparisons in each of these settings separately. The prior unlearning methods that do not adapt to the setting where there is no access to $\mathcal{D_{\text{R}}}$ Warnecke et al. (2021); Liu et al. (2024) are excluded for the presented results in that setting.

### v3

@@ 1 Introduction
The goal of machine unlearning is to remove the influence of a subset of the training dataset for a model that has been trained on that dataset (vatter2023evolution). The necessity for these methods has been determined by privacy regulations such as the European Union’s General Data Protection Act and the California Consumer Privacy Act. Despite early efforts on proposing “exact” solutions to this problem (cao2015towards; bourtoule2021machine), the community has favored “approximate” solutions due to their ability to preserve the original model’s accuracy while being more computationally efficient (chen2023boundary; liu2024model; fan2023salun).
@@
As we will show in § 6, AMUN outperforms prior state-of-the-art (SOTA) unlearning methods (fan2023salun) in unlearning random subsets of the training data from a trained classification model and closes the gap with the retrained models, even when there is no access to the samples in $\mathcal{D}-\mathcal{D_{\text{F}}}$ during the unlearning procedure. The code for reproducing the results can be found in https://github.com/Ali-E/AMUN.
@@ 1.1 Related Work
Early works in machine unlearning focused on exact solutions cao2015towards; bourtoule2021machine; those ideas were adapted to unlearning in other domains such as graph neural networks chen2022graph and recommendation systems chen2022recommendation. The extensive computational cost and utility loss resulted in the design of approximate methods. An example is the work of ginart2019making, who provide a definition of unlearning based on differential privacy. Works that followed sought solutions to satisfy those probabilistic guarantees (ginart2019making; gupta2021adaptive; neel2021descent; ullah2021machine; sekhari2021remember). However, the methods that satisfy these guarantees were only applied to simple models, such as $k$-means ginart2019making , linear and logistic regression guo2019certified; izzo2021approximate, convex-optimization problems neel2021descent, or graph neural networks with no non-linearities chien2022efficient. Additional research was carried out to design more scalable approximate methods, those that can be applied to the models that are used in practice, including large neural networks golatkar2020eternal; warnecke2021machine; izzo2021approximate; thudi2022unrolling; chen2023boundary; liu2024model; fan2023salun. However, these approximate methods do not come with theoretical guarantees; their effectiveness are evaluated using membership inference attacks (MIAs). MIAs aim to determine whether a specific data sample was used in the training set of a trained model (shokri2017membership; yeom2018privacy; song2019privacy; hu2022membership; carlini2022membership; zarifzadeh2024low), and is a common evaluation metric (liu2024model; fan2023salun). s For further discussion on related works see Appendix B.
@@ 3.1 A Guiding Observation
Results: Figure 3 in Appendix E.1, shows the confidence scores (see § E.1 for details) for a ResNet-18 he2016deep model that has been retrained on $\mathcal{D}-\mathcal{D_{\text{F}}}$, where $\mathcal{D}$ is the training set of CIFAR-10 alex2009learning and the size of $\mathcal{D_{\text{F}}}$ (randomly chosen from $\mathcal{D}$) is $10\%$ and $50\%$ of the size of $\mathcal{D}$ (the first and second sub-figures, respectively).
@@
This basic observation has either been overlooked by the prior research on approximate machine unlearning or has been treated incorrectly. To make the unlearned models more similar to $\mathcal{F_{\text{R}}}$, prior methods have focused on degrading the model’s performance on $\mathcal{D_{\text{F}}}$ directly by either (a) some variation of fine-tuning on $\mathcal{D_{\text{R}}}$ (warnecke2021machine; liu2024model), (b) choosing wrong labels for samples in $\mathcal{D_{\text{F}}}$ and fine-tuning the model (golatkar2020eternal; chen2023boundary; fan2023salun), or (c) directly maximizing the loss with respect to the samples in $\mathcal{D_{\text{F}}}$ (thudi2022unrolling). Using the wrong labels for the samples of $\mathcal{D_{\text{F}}}$ or maximizing the loss on them make these methods very unstable and prone to catastrophic forgetting (zhang2024negative) because these samples belong to the correct distribution of the data and we cannot force a model to perform wrongly on a portion of the dataset while preserving it’s test accuracy. For these methods, it is important to use a small enough learning rate along with early stopping to prevent compromising the model’s performance while seeking worse prediction confidence values on the samples in $\mathcal{D_{\text{F}}}$. Also, most of these methods require access to the set of remaining samples to use it for preventing a total loss of the model’s performance (e.g., by continuing to optimize the model on $\mathcal{D_{\text{R}}}$(golatkar2020eternal; liu2024model).
@@ 3.2 Fine-tuning on Adversarial Examples
Now, what happens if we insert one adversarial example $(x_{adv},y_{adv})$ that corresponds to the sample $(x,y)$ into $\mathcal{D}$ and make an augmented dataset $\mathcal{D}^{\prime}$ for fine-tuning? Even before fine-tuning starts, the model makes the correct prediction on that (adversarial) example (by predicting the wrong label $y_{adv}$!), but its confidence might not be as high as the samples in $\mathcal{D}$, on which the model has been trained on. Proceeding with fine-tuning of the model on the augmented dataset increases its confidence on $x_{adv}$ while making the same wrong prediction $y_{adv}$. However, this fine-tuning does not change the model’s performance because the newly added sample $(x_{adv},y_{adv})$ does not contradict the distribution learned by the model. Since $(x,y)\in\mathcal{D}^{\prime}$, and $x$ and $x_{adv}$ are very close to one another (e.g., very similar images) while having different labels, optimizing the model has to change its decision boundary in that region of the input space to reach small loss for both of these samples. As a result of this balance, the model tends to decrease its confidence on the original sample compared to the model that was solely trained on $\mathcal{D}$ because there was no opposing components for its optimization on $\mathcal{D}$. Note that $\|x-x_{adv}\|\leq\epsilon$, where $\epsilon$ is often much smaller than the distance of any pairs of samples in $\mathcal{D}$. This helps to localize this change in the decision boundary during fine-tuning, and prevent changes to models’ behavior in other regions of the input space (liang2023towards). In the following we elaborate on our empirical observations that verify these changes.
@@ 4 Adversarial Machine UNlearning (AMUN)
Formally, AMUN uses Algorithm 1 to find an adversarial example for any sample in $(x,y)\in\mathcal{D_{\text{F}}}$. This algorithm uses a given untargeted adversarial algorithm $\mathcal{A}_{\mathcal{F}}$, that finds the solution to Definition˜2.1, for finding an adversarial example $x_{adv}$. To make sure $\epsilon$ is as small as possible, Algorithm 1 starts with a small $\epsilon$ and runs the attack $\mathcal{A}_{\mathcal{F}}$; if an adversarial algorithm is not found within that radius, it runs $\mathcal{A}_{\mathcal{F}}$ with a larger $\epsilon$. It continues to perform $\mathcal{A}_{\mathcal{F}}$ with incrementally increased $\epsilon$ values until it finds an adversarial example; it then adds it to $\mathcal{D_{\text{A}}}$. The algorithm stops once it finds adversarial examples for all the samples in $\mathcal{D_{\text{F}}}$.
The reason behind minimizing the distance of $\epsilon$ for each sample is to localize the changes to the decision boundary of the model as much as possible; this prevents changing the model’s behavior on other parts of the input space. For our experiments, we use PGD-50 (madry2017towards) with $\ell_{2}$ norm bound as $\mathcal{A}_{\mathcal{F}}$. We set the step size of the gradient ascent in the attack to $0.1\times\epsilon$, which changes with the $\epsilon$ value. More details regarding the implementations of AMUN and prior unlearning methods and tuning their hyper-parameters can be found in Appendix C. Also, in Appendix E.5, we will show how using weaker attacks, such as Fast Gradient Sign Method (FGSM) (goodfellow2014explaining), might lead to lower performance of AMUN.
@@ 4.1 Influencing Factors
In this section, we derive an upper-bound on the $2$-norm of the difference of the parameters of the unlearned model and the retrained model (which are gold-standard for unlearning) that illuminated the influencing factors in the effectiveness of AMUN. To prove this theorem, we make assumptions that are common in the certified unlearning literature. The proof is given in Appendix˜A.
@@
Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=\{1,\dots,N\}}$ be a dataset of $N$ samples and without loss of generality let $(x_{n},y_{n})$ (henceforth represented as $(x,y)$ for brevity) be the sample that needs to be forgotten and $(x^{\prime},y^{\prime})$ be its corresponding adversarial example used by AMUN such that $\|x-x^{\prime}\|_{2}=\delta$. Let $\hat{\mathcal{R}}(w)$ represent the (unnormalized) empirical loss on $\mathcal{D}^{\prime}=\mathcal{D}\cup\{(x^{\prime},y^{\prime})\}$ for a model $f$ that is parameterized with $w$. We assume that $f$ is $L$-Lipschitz with respect to the inputs and $\hat{\mathcal{R}}$ is $\beta$-smooth and convex with respect to the parameters. Let $\theta_{o}$ represent the parameters corresponding to the model originally trained on $\mathcal{D}$ and $\theta_{u}$ be the parameters derived when the model is trained on $\mathcal{D}-\{(x,y)\}$. We also assume that both the original and retrained models achieve near-0 loss on their training sets. After AMUN performs fine-tuning on $\mathcal{D}^{\prime}$ using one step of gradient descent with a learning rate of $\frac{1}{\beta}$ to derive parameters $\theta^{\prime}$, we get the following upper-bound for the distance of the unlearned model and the model retrained on $\mathcal{D}-\{(x,y)\}$ (gold standard of unlearning):
@@
Note that the first two implications rely on the strength of the adversarial example in addition to being close to the original sample. The second bullet, which relies on the transferability of adversarial examples, has been shown to improve as the Lipschitz constant decreases (ebrahimpourtraining). The third bullet point is a natural implication which also holds for other unlearning methods that rely on the fine-tuning of the model. The fourth bullet point is not relevant to the unlearning method and instead relies on the fact that the retrained model should have good generalizability to unseen samples; it implies that as the size of $\mathcal{D_{\text{F}}}$ increases (i.e., $|\mathcal{D_{\text{R}}}|$ decreases) and the performance of the retrained model decreases, the effectiveness of the unlearning model also decreases. This is also intuitively expected in the unlearning process. Hence, the proved theorem also justifies our earlier intuitions about the need for good quality adversarial examples that are as close as possible to the original samples (which is the goal of Algorithm 1).
@@ 5.1 Baseline Methods
We compare AMUN with FT (warnecke2021machine), RL (golatkar2020eternal), GA (thudi2022unrolling), BS (chen2023boundary), $l_{1}$-Sparse (liu2024model), and SalUn (fan2023salun). We also combine the weight saliency idea for masking the model parameters to limit the changes to the parameters during fine-tuning with AMUN and present its results as AMUN+SalUn (see Appendix D for more details). We use the same hyper-parameter tuning reported by prior works. For further details, see Appendix C.
@@ 5.2 Evaluation Metrics
The metic used by recent works in unlearning to evaluate the unlearning methods (liu2024model; fan2023salun) considers the models retrained on $\mathcal{D_{\text{R}}}$ as the goal standard for comparison. They compute the following four values for both the retrained models and the models unlearned using approximate methods:
@@
The MIAs used in the recent unlearning methods by liu2024model; fan2023salun are based on the methods introduced by yeom2018privacy; song2019privacy. Although these MIAs have been useful for basic comparisons, recent SOTA MIAs significantly outperform their earlier counterparts, albeit with an increase in complexity and computation cost. To perform a comprehensive comparison of the effectiveness of the unlearning methods, we utilized a SOTA MIA called RMIA (zarifzadeh2024low), in addition to using the MIAs from prior works. In RMIA, the area under the ROC curve (AUC) of the MIA scores for predicting the training samples from the unseen samples is reported. Recall that in machine unlearning, the samples are split to three sets: $\mathcal{D_{\text{R}}}$, $\mathcal{D_{\text{F}}}$, and $\mathcal{D_{\text{T}}}$. For an unlearning method to be effective, as discussed in § 3.1, we expect the AUC of RMIA for distinguishing the samples in $\mathcal{D_{\text{F}}}$ from the ones in $\mathcal{D_{\text{T}}}$ to be the same as random guessing ($50\%$ assuming balanced data). As shown in Table 1, this expectation holds for the models retrained on $\mathcal{D_{\text{R}}}$.
@@ 5.3 Unlearning Settings
Another important factor missing in the comparisons of the unlearning methods in prior works is the possibility of access to $\mathcal{D_{\text{R}}}$. So, for our experiments we consider two settings, one with access to $\mathcal{D_{\text{R}}}$ and one with access to only $\mathcal{D_{\text{F}}}$. We adapt each of the unlearning methods to both of these settings, and perform the comparisons in each of these settings separately. The prior unlearning methods that do not adapt to the setting where there is no access to $\mathcal{D_{\text{R}}}$ warnecke2021machine; liu2024model are excluded for the presented results in that setting.